b) A government bond that originally cost $5000 with a yield of 6% (simple interest) has 5 years left before redemption.


i. Determine its present value if the prevailing rate of interest is 10%. Briefly explain the steps in your own words.


ii. Is it worth purchasing this bond? Provide your own reasoning.


1
Expert's answer
2021-05-26T15:55:38-0400

b) i. we will find it by the formula

"PV=\frac{C}{(1+r)^1}+\frac{C}{(1+r)^2}+\frac{C}{(1+r)^3}+\frac{C}{(1+r)^4}+\frac{C}{(1+r)^5}+\frac{N}{(1+r)^5}"

N=5000

"C=5000\times0.06=300"

"PV=\frac{300}{(1+0.1)^1}+\frac{300}{(1+0.1)^2}+\frac{300}{(1+0.1)^3}+\frac{300}{(1+0.1)^4}+\frac{300}{(1+0.1)^5}+\frac{6000}{(1+0.1)^5}=4862.76"

ii. This bond is sold at a discount, so it is a profitable investment, so the investor will pay less to get more in the future. Also, all other things being equal, the purchase will pay off, since the price is less than the nominal value.